The health care provider (HCP) orders 10 mg/kg/day. The patient weight is 12 kg. The medication...

  • The health care provider (HCP) orders 10 mg/kg/day. The patient weight is 12 kg. The medication label reads 50 mg/mL. What is the divided dose to be administered?
Answer #1

patient weight = 12 Kg

the dose per day = 10mg/kg / day

the dose per day for the child = 10 x 12 = 120mg/day

There fore if the medicine is to begiven three times a day, the divided dose = 120/3 = 40 mg

the medication label says each ml contains 50mg, then how many ml has to be administered per dose?

= 40x1 /50 = 0.8 ml